OpenStudy (anonymous):

how can I show my work on -2n+5=-15 (solve the equation)

OpenStudy (saifoo.khan):

hi rose!

OpenStudy (anonymous):

hey whats up? will you help me a couple mins?

OpenStudy (saifoo.khan):

sure.. Im here! :)

OpenStudy (anonymous):

5t-1=-11

OpenStudy (saifoo.khan):

5t = -11 + 1 5t = 10 t = 10/5 t = 2

OpenStudy (anonymous):

-4=-3p+14

OpenStudy (anonymous):

I'm surprised you didn't ask here how old I was , lol

OpenStudy (anonymous):

how old?

OpenStudy (saifoo.khan):

-2p + 14 = -4 -2p = -4 -11 -2p = -15 p = 15/2

OpenStudy (saifoo.khan):

he is 49

OpenStudy (anonymous):

Wrong! lol

OpenStudy (anonymous):

lol

OpenStudy (saifoo.khan):

94?

OpenStudy (anonymous):

Don't tell him

OpenStudy (anonymous):

Saifoo, you're the one that's old. You're not funny either.

OpenStudy (anonymous):

You're just a little joker

OpenStudy (saifoo.khan):

Dont tell your case plz Zero.

OpenStudy (anonymous):

hey number 3 is not -6

OpenStudy (anonymous):

Are you doing another online test again?

OpenStudy (saifoo.khan):

where?

OpenStudy (anonymous):

the last one u just did is not -6?

OpenStudy (saifoo.khan):

-2p + 14 = -4

OpenStudy (saifoo.khan):

this ome?

OpenStudy (anonymous):

Hero I have a final on wends trying to get theseright.

OpenStudy (anonymous):

-4=-3p+14

OpenStudy (saifoo.khan):

oh. okay!

OpenStudy (saifoo.khan):

-3p + 14 = -4 -3p = -4 -14 -3p = -18 3p = 18 p = 18/3 p = 6

OpenStudy (anonymous):

why did u give me a diff. answer you better be right this is homework. im depending on u.

OpenStudy (anonymous):

lol, he's been a little rusty today on answers, lol

OpenStudy (anonymous):

hes working overtime.

OpenStudy (anonymous):

and not getting paid for it either, lol

OpenStudy (anonymous):

well his brains are and its good he can help us..... so whats up kuan with the mishap

OpenStudy (anonymous):

What are these...a bunch of review problems or something?

OpenStudy (anonymous):

yeah how old r u by the way?

OpenStudy (anonymous):

I thought I told you already....Aren't you the girl from Florida?

OpenStudy (anonymous):

oh you live in saudia?

OpenStudy (anonymous):

I'm from Chicago...lol

OpenStudy (anonymous):

How did you know i lived in fla. No you never told me your age....

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

The only people I am familar with is kuan and this man with a masters degree in statistics and hes pure genius

OpenStudy (anonymous):

Saifoo thinks you know my age

OpenStudy (anonymous):

he lives in saudia neways would u help me i have to finish

OpenStudy (anonymous):

well I dont

OpenStudy (anonymous):

Well that's good

OpenStudy (anonymous):

How many problems do you have? Don't lie

OpenStudy (anonymous):

18

OpenStudy (anonymous):

why because u lie so u thnk im going to lie?

OpenStudy (anonymous):

Wait, who said I lied, lol

OpenStudy (anonymous):

http://www.twiddla.com/solved

OpenStudy (anonymous):

that site dont work

OpenStudy (anonymous):

Try again, lol

OpenStudy (anonymous):

What do you mean it doesn't work....lol

OpenStudy (anonymous):

I tryed and its just clicking back and forth like theres a issue

OpenStudy (anonymous):

Oh well, I'm sleepy now...

OpenStudy (anonymous):

I'm sure saifoo will help you with the rest

OpenStudy (anonymous):

I'm surprised he hadn't came yet...

OpenStudy (anonymous):

Were you about to say something?

jhonyy9 (jhonyy9):

-2n=-20 n=-20/-2 =10 n=10